4) Anders Inc. uses the weighted-average method in its process costing system. The following data concern the operations of the companys first processing department for a recent month. Work in process, beginning Units in process Percent complete with respect to materials Percent complete with respect to conversion Costs in the beginning inventory Materials cost Conversion cost 800 70% 10% S 1,736 S 2,288 Units started into production during the month Units completed and transferred out 18,000 18,400 Costs added to production during the month Materials cost Conversion cost S 59,908 S 520,696 Work in process, ending Units in process Percent complete with respect to materials Percent complete with respect to conversion 400 70% 20% Required: a. Determine the equivalent units of production. b. Determine the costs per equivalent unit c. Determine the cost of ending work in process inventory d. Determine the cost of the units transferred to the next department
Answer #1

a) Equivalent unit of production

Physical Unit % EUP-Material % EUP-Conversion
Unit transferred out 18400 100% 18400 100% 18400
Ending work in process unit 400 70% 280 20% 80
Equivalent unit of production 18680 18480

b) Calculate cost per equivalent unit

Material Conversion
Beginning WIP 1736 2288
Cost added 59908 520696
Total Cost 61644 522984
Equivalent unit 18680 18480
Cost per equivalent unit 3.3 28.3

c) Cost of ending work in process = (280*3.3+80*28.3) = $3188

d) Cost of unit transferred out = (3.3+28.3)*18400 = $581440